OpenStudy (anonymous):

A triangle has the following side lengths: 3 3/5 inches and 4 5/8 inches. The total distance around the triangle, perimeter, is 15 1/6 inches. What is the length of the missing side?

OpenStudy (aaronandyson):

Missing length = perimeter - side 1 - side 2 15 1/6 - 3 3/5 - 4 5/8 = 301/120 - 75/120 - 65/120 = 161/120 = 1 41/120 inches

OpenStudy (anonymous):

Thank you.

OpenStudy (aaronandyson):

np :)